Session: Oct 2012 (session 3)
Wavelength: 5 cm
To give feedback on this experiment, click
(scalambra at Sun Nov 04 19:00:59 2012) Success:
Yebes (Daniel Sastre at Sun Nov 04 19:09:51 2012) Minor failures: WIND
deep fog and rain during the entire observation
(Michael Lindqvist at Sun Nov 04 19:42:35 2012) Success:
(K. Borkowski at Mon Nov 05 07:35:30 2012) Minor failures: RECO
Part of scan 14 (that started at 14:49 UT) not recorded due to problems with the disk-pack (Mark5A issued repeatedly the warning: "Recording throttled (can't keep up)" with record pointer not incrementing, although the pack had still much free space). Otherwise no known problems.
Effelsberg (GEORGI at Mon Nov 05 09:33:34 2012) Success:
(Geert Kuper at Tue Nov 06 11:35:09 2012) Minor failures:
12:19 UT - 14:30 Low TP level due to backend attenuator steering problem
14:15 UT - 14:30 UT no signal due to reboot backend
14:30 UT - 15:52 UT Formatter sync error gt 3 microsec
(Geert Kuper at Tue Nov 06 11:57:52 2012) Success:
Additional info for es069b:
UT 16:48 - 17:18 no data due to backend reset and test.
(Alastair Gunn at Mon Nov 12 16:11:06 2012) Success:
No known problems.